The force on bullet is given by the formula F=580-(1.8x10^5)t over the time interval t=0 to t=3x10^-3s. F is newtons. Sketch a graph F vs t for t=0 to t=3x10^-3


1
Expert's answer
2021-01-28T09:57:52-0500

Explanations & Calculations


  • Equation is given as "\small F=-(1.8\times 10^5)t+580" which is a linear function of time & it is in the form "\small y=-mx+c"
  • To plot a linear function coordinates of just 2 different points are needed.
  • Therefore,

@ t=0: F=580 and @ t="\small 3\times 10^{-3}" : F=40

Additionally, @F=0 : t= "\small 3.2222\times 10^{-3}"

  • Then graph can be plotted to be like this